OpenStudy (anonymous):

A sum of money double in 20 years on simple interest . it will triple at the same rate in _?

OpenStudy (anonymous):

no thr no option of 30

OpenStudy (anonymous):

40 60 50 80

OpenStudy (phi):

formula for simple interest is I = P r t when I = P (the amount of interest equals the principal) you have doubled your money so P= P r 20 r= 1/20 now solve for when I= 2P

OpenStudy (anonymous):

higs can you solve it

OpenStudy (anonymous):

phi iknow the formula but dont get the concept of that q

OpenStudy (shane_b):

I think you should look closer at Phi's answer :)

OpenStudy (shane_b):

Since it doubles in 20 years that means every 20 years you gain P. It will take 40 years to gain P two times. When you add this to the original P you end up with triple P.

OpenStudy (phi):

you said "i know the formula but dont get the concept of that q" Well, first think of this. You have some money. Call the amount P. How much more do you need to double the amount of money you have? So if you get interest, you want that interest to be the amount needed to double your money. Can you use the formula to find the rate of interest. That is the first part of the problem. Second part of the problem. You start with P amount of money. How much more do you need to triple it? What is triple P? Yes it is 3P. that means you need 2P more. use the formula again, with I= 2P, r= 1/20 (found from the first part). use algebra to find the time.

OpenStudy (anonymous):

i got it now thanks

OpenStudy (anonymous):

its 40